Project

General

Profile

Actions

Ops/Comm Scheduling meeting : 12/9/2017

Location : CR21 at 14:00 pm

Phone access : phonebridge number
UK: 08082380274
US: 631.621.5253
Access code: 1779339517

  • Actions
    • Give LH2 documentation to Steve to add to official documentation list (Craig, Josef): completed draft exists but further changes expected; stands
    • Arrange a geometry survey for September just before running starts (Colin/ChrisR) : Tony Millington generally available, fix date closer to the time; stands
  • DC update
  • Pulser run summary (MOM)
    • EMR DAQ arming problem
  • Hydrogen system plan (JBoehm)
    • Documentation required for CR and for training on Monday 18th
  • Status of alarm clearance (DC/Ajit)
    • Report from Ajit (who can't make this meeting)
      • Ran through current alarms with Paolo and identified which were due to hardware issues, software issues and those that needed limits to be changed. I am in the process of fixing the software related issues.
      • Paolo will follow up these issues with the system owners.
      • The system temperature alarms on the ioc machines will have their limits increased to allow for the fact that some are running hot. High limit will be increased to 48 and HIHI to 52 on machines that are hot. These are still less than the peak temperatures recorded in July.
  • Update on tracker readout reconfiguration (or not) (Melissa)
  • Update on Q8 (Henry)
  • Update on DS Shield pressure (Henry)
  • File Compactor vs DAQ (Henry/Durga)
  • Update on SS QPS connector issue (Craig)
  • Operations and Procedures?
  • Operations startup
    • Thursday beam run plan
    • BLOC signup
    • Monday Sep 18th is the first training day. Starts at 10:30 am in the MICE conference room in R1
  • Update on Stakeholder Transfer of Engagement And Knowledge (CWhyte)
    Concerns raised regarding Cost-Benefit :(
  • AOB

Updated by Boyd, Steven almost 5 years ago · 10 revisions